# Yang-Lee Quantum Criticality in Various Dimensions

### v1

@@ Abstract
The Yang-Lee universality class arises when imaginary magnetic field is tuned to its critical value in the paramagnetic phase of the $d<6$ Ising model. In $d=2$, this non-unitary Conformal Field Theory (CFT) is exactly solvable via the $M(2,5)$ minimal model. As found long ago by von Gehlen using Exact Diagonalizations, the corresponding real-time, quantum critical behavior arises in the periodic Ising spin chain when the imaginary longitudinal magnetic field is tuned to its critical value from below. Even though the Hamiltonian is not Hermitian, the energy levels are real due to the $\mathcal{PT}$ symmetry. In this paper, we explore the analogous quantum critical behavior in higher dimensional non-Hermitian Hamiltonians on regularized spheres $S^{d-1}$. For $d=3$, we use the recently invented, powerful fuzzy sphere method, as well as discretization by the platonic solids cube, icosahedron and dodecaherdron. The low-lying energy levels and structure constants we find are in agreement with expectations from the conformal symmetry. The energy levels are in good quantitative agreement with the high-temperature expansions and with Padè extrapolations of the $6-\epsilon$ expansions in Fisher’s $i\phi^{3}$ Euclidean field theory for the Yang-Lee criticality. In the course of this work, we clarify some aspects of matching between operators in this field theory and quasiprimary fields in the $M(2,5)$ minimal model. For $d=4$, we obtain new results by replacing the $S^{3}$ with the self-dual polytope called the $24$-cell.
@@ 1 Introduction and Summary
The Ising model is the most fundamental and widely explored statistical model, which has a broad range of applications across sciences. Its one-dimensional version was solved by Ernst Ising in 1925 and shown not to possess a finite-temperature phase transition [1]. In all dimensions $d>1$, however, it has a line of first-order transitions $T<T_{c}$, which terminates at the critical point where the $Z_{2}$ symmetry is restored. The properties of this second-order phase transition have been studied using many different methods, analytical and numerical. For $d>4$, the critical exponents of the Ising model are given exactly by Landau’s mean field theory, but for $d<4$ this is no longer the case. Determining the dimension dependence of the critical exponents is a fascinating problem, which continues to attract a great deal of attention.
@@
In $d=2$, the exact Ising critical exponents have been known for many years thanks to the Onsager solution [2]. The critical 2D Ising model corresponds to the simplest unitary Belavin-Polyakov-Zamolodchikov (BPZ) conformal minimal model $M(3,4)$ [3]. More generally, the critical $d$-dimensional Ising model is described by the continuum theory of a massless scalar field $\phi$ with the interaction term $\sim\phi^{4}$. The problem of finding the critical exponents is then mapped to calculation of the scaling dimensions of operators $\phi$, $\phi^{2}$, $\phi^{4}$, etc. In dimensions slightly below $4$, the IR fixed point is weakly coupled, and the Wilson-Fisher $4-\epsilon$ expansion [4] provides a useful method for approximating the $d$ dependence of operator dimensions. This has led to estimates of the 3D Ising critical exponents that are in very good agreement with experiments and Monte Carlo simulations. A significant improvement in accuracy was provided by the numerical conformal bootstrap [5, 6], which is based on the conformal symmetry of critical fluctuations [7] (for reviews, see [8, 9, 10, 11]). Furthermore, the numerical bootstrap method has been applied to the operator scaling dimensions in the entire range $42\leq d<$ 4 [12], producing excellent agreement with resummed $4-\epsilon$ expansions for the $\phi^{4}$ theory.
@@
More recently, new methods for studying the 3D critical phenomena have emerged [13, 14, 15, 16, 17]. They rely on the critical behavior of the quantum theory in $2+1$ dimensions, which is analogous to the well-known periodic quantum spin chains in $1+1$ dimensions. In the continuum limit, the latter give the energy levels on a circle which, via the 2D Conformal Field Theory state-operator map determine the scaling dimensions of the primary operators and their descendants. To study a 3D CFT in a similar way, one must define the quantum Hamiltonian system on an appropriately regularized sphere $S^{2}$. In an important development [13], the fuzzy sphere regularization [18, 19] was shown to work very nicely for the critical 3D Ising model, giving the low-energy states that organize into the expected conformal multiplets (in fact, replacing the $S^{2}$ by an icosahedron works quite well too [16], even though this method does not readily allow for extrapolation to the thermodynamic limit). Besides the scaling dimensions and OPE coefficients in a 3D CFT, the fuzzy sphere approach provides access to the sphere free energy $F$, which characterizes the number of degrees of freedom [20, 21, 22]. This quantity can be extracted from the entanglement entropy across the equator of $S^{2}$, and the numerical result for $F$ in the Ising model [15] is in excellent agreement with the extrapolation of the $4-\epsilon$ expansion of $F$ [23].
@@
A fascinating feature of the $d$-dimensional Ising model is the appearance of zeros of the partition function $Z$ at imaginary values of the magnetic field [24, 25]. In the paramagnetic phase $T>T_{c}$, these Yang-Lee (YL) zeros become dense near the edge value $h_{\rm crit}(T)$. As shown by Kortman and Griffiths [26], the density of zeros has the scaling form $\rho(h)\sim(h-h_{\rm crit}(T))^{\sigma}$, and $\sigma$ is negative in $d=2$ but positive in $d=3$. In an important further development, Michael Fisher [27] realized that this critical behavior is described by the Euclidean field theory with the imaginary potential $\sim i\phi^{3}$. In this continuum approach, the critical exponent is $\sigma=\frac{\Delta_{\phi}}{d-\Delta_{\phi}}$, where $\Delta_{\phi}$ is the dimension of the real scalar field $\phi$. The field theory description of the YL universality class implies that its upper critical dimension is $6$. The lower critical dimension is $1$, and the exact solution there gives $\sigma=-\frac{1}{2}$ [24, 25]. Therefore, an interesting problem is to determine the $d$ dependence of the critical exponents in the range $61<d<$ 6. One approach, analogous to the Wilson-Fisher $4-\epsilon$ expansion for the Ising universality class, is to develop the $6-\epsilon$ expansion for the massless $i\phi^{3}$ field theory. This approach, in conjunction with the exact result in $d=1$, was used by Fisher [27] to estimate the $d$-dependence of $\sigma$ for the YL universality class. Soon after the advent of the BPZ minimal models [3], Cardy [28] realized that the critical YL model is also exactly solvable in $d=2$, where it corresponds to the simplest non-unitary minimal model $M(2,5)$ (for a recent review, see [29]). The field $\phi$ corresponds to the Virasoro primary $\Phi_{1,2}$ of holomorphic dimension $h_{1,2}=-\frac{1}{5}$. This means that $\Delta_{\phi}=-\frac{2}{5}$ and $\sigma=-\frac{1}{6}$ in $d=2$.
@@
In the range $62<d<$ 6, no similar exact results are known. In this paper, we will use a combination of $6-\epsilon$ expansions and numerical diagonalizations to approximately determine the scaling dimensions of various operators. The $6-\epsilon$ expansions are now available to $O(\epsilon^{5})$ [30]. Their resummations have been extrapolated to $d=2$, yielding good agreement with $M(2,5)$. However, for the higher dimension operators this method can exhibit significant uncertainties. The numerical boostrap approach is complicated by the fact that the YL model is non-unitary, so that the standard methods based on positivity cannot be used. A different bootstrap approach to the YL model was used in [31, 32, 33], but it does not appear to be as precise as the numerical bootstrap based on positivity. We will instead follow the successes of the fuzzy sphere [13] and the icosahedron[16] [^1] Icosahedron has also been used to discretize classical Ising model in [34, 35] approaches to the 3D Ising model, adapting them to the YL case where a non-Hermitian Hamiltonian has to be tuned to quantum criticality. We will proceed analogously to the quantum spin chain approach which has proven to be successful for describing the 2D YL model. Here an imaginary longitudinal magnetic field $h^{z}$ was introduced and tuned to a critical value from below [36, 37, 38]. Even though the spin chain Hamiltonian is complex, the energy levels are real for $h^{z}<h^{z}_{\rm crit}$ due to the $\mathcal{PT}$ symmetry [38].
@@
This paper is organized as follows. In section 2 we review the minimal model $M(2,5)$ and its field theory description, discussing the correspondence between the two. In section 3 we review the quantum spin chain with an imaginary longitudinal magnetic field [39, 36, 37, 38]. We carry out its numerical diagonalizations and show that the appropriate criteria for quantum criticality, analogous to those used recently [40] for a precise study of the Ising transition in the Schwinger model, lead to excellent agreement with the exact results from the $M(2,5)$ minimal model. In section 4 we study a quantum spin model with an imaginary longitudinal magnetic field, putting it on the platonic solids icosahedron and dodecahedron, which are viewed as regularized versions of $S^{2}$. We obtain good agreement with the expected YL criticality. In section 5, we carry out analogous calculations on the fuzzy sphere and extrapolate them to the thermodynamic limit. Our fuzzy sphere result for the scaling dimension of $\phi$, $\Delta_{\phi}=0.214(2)$ is in excellent agreement with the high-temperature expansion result $0.214(6)$ from [41] and is also very close to our two-sided Padé estimate $\approx 0.218$. In section 6, we obtain first results for the quantum criticality of the $d=4$ YL model by replacing the $S^{3}$ by the $24$-cell, which is a selfdual polytope.
@@ 2 Minimal model $M(2,5)$ and its field theory description
The lowest level null vectors in the $M(2,5)$ model are [45]
$L_{-1}|I\rangle,\quad\Big{(}L_{-2}^{2}+\frac{3}{5}L_{-4}\Big{)}|I\rangle,\quad\Big{(}L_{-2}-\frac{5}{2}L_{-1}^{2}\Big{)}|\phi\rangle,\quad\Big{(}L_{-3}-\frac{10}{3}L_{-1}L_{-2}+\frac{25}{36}L_{-1}^{3}\Big{)}|\phi\rangle\,.$ (2.3)
All the operators (primary and descendant) up to level $4$ are exhibited in Figure 1, where $T=L_{-2}|I\rangle$ is the holomorphic part of the stress-energy tensor and
$Q\phi=(L_{-4}-625/624L_{-1}^{4})|\phi\rangle\ ,\qquad\bar{Q}\phi=(\bar{L}_{-4}-625/624\bar{L}_{-1}^{4})|\phi\rangle\ .$ (2.4)
denote the spin $4$ quasi-primary state [46]. A related spin 00 quasiprimary is $\Xi=Q\bar{Q}|\phi\rangle$.
Using the operator $Q_{6}=L_{-6}+...$, we can also construct a spin $6$ quasiprimary state:
$Q_{6}|\phi\rangle\ ,\qquad\bar{Q}_{6}|\phi\rangle\ \ ,$ (2.5)
@@
The operator $\phi^{2}$ is a descendant by the equation of motion. Therefore, the next spin-zero quasiprimary field is $\phi^{3}$. Its scaling dimension is [48, 49, 50, 30]
@@
The field theory also contains quasiprimary operators with an even number of derivatives, which have schematic form $\phi\partial_{\mu_{1}}\partial_{\mu_{2}}\ldots\partial_{\mu_{2n}}\phi+\ldots$. Projecting onto the traceless part of such an operator, we find the leading twist operators of spin $2n$ (they form the leading Regge trajectory). For $n=1$ this operator is the conserved stress-energy tensor, but for $n>1$ they are not conserved. The first such operator is the traceless rank $4$ tensor
@@
Table 2: Two-sided Padé approximations of the scaling dimensions of lowest quasiprimary operators in $d=3,4,5$. All of them are below the unitarity bounds.
@@ 3 Spin chain for 2D Yang-Lee model
Figure 4: The qualitative phase diagram of the quantum Ising model with transverse magnetic field $h_{x}$ and imaginary longitudinal magnetic field $h_{z}$. The $h_{z}=0$ horizontal line is the $\mathbb{Z}_{2}$ symmetric line. The Ising critical point, which is located at this line, is shown by the green dot. To the right of this point, the Ising model is in the paramagnetic phase, also known as the high temperature phase, shown by the black section of the $\mathbb{Z}_{2}$ symmetric line. The paramagnetic phase with a small $h_{z}$ deformation still preserves $\mathcal{PT}$ symmetry, shown by the white region. When $h_{z}$ passes its critical value, the $\mathcal{PT}$ symmetry is spontaneously broken, shown by the blue region. The critical coupling $h_{z}$ depends on $h_{x}$ and therefore we have a line of critical points corresponding to the YL criticality, shown as the blue solid line. We study the Ising criticality to YL criticality flow by moving along the red dashed arrows between points a, b and c.
@@
where, in the numerical spectrum, this ratio can be written using the energy levels of distinct spin sectors, namely, $E^{(\textbf{s})}_{n}$ with $n=0,1,2,\dots$ represents the $n$th energy level in spin sector $\textbf{s}=0,1,2,\dots$, so
@@
to be exactly $r_{T}=3$. Numerically this ratio can be obtained using the following energy levels in different irreps of $O$ and $I_{h}$:
@@
where we used that $\textbf{5}_{O(3)}$ splits as $\textbf{2}+\textbf{3}^{\prime}$ under the cubic symmetry group $O$, whereas it does not split for Icosahedral group $I_{h}$, and $\textbf{1}_{O(3)}$ and $\textbf{3}_{O(3)}$ do not split for $O$ and $I_{h}$ and corresponds to irrep 1 and 3. Here we list values $ih_{z}^{*}$, where $r_{T}=3$ in the Table 7. They are slightly away from the merger points listed in the previous Table 6.
@@
$\textrm{Cube}:\quad\Delta_{\phi}=3\frac{E_{1}^{(\textbf{1})}-E_{0}^{(\textbf{1})}}{E_{1}^{(\textbf{3}^{\prime})}-E_{0}^{(\textbf{1})}}\,,\qquad\textrm{Ico/Dode}:\quad\Delta_{\phi}=3\frac{E_{1}^{(\textbf{1})}-E_{0}^{(\textbf{1})}}{E_{1}^{(\textbf{5})}-E_{0}^{(\textbf{1})}}\,,$ (4.4)

### v2

@@ Abstract
The Yang-Lee universality class arises when imaginary magnetic field is tuned to its critical value in the paramagnetic phase of the $d<6$ Ising model. In $d=2$, this non-unitary Conformal Field Theory (CFT) is exactly solvable via the $M(2,5)$ minimal model. As found long ago by von Gehlen using Exact Diagonalization, the corresponding real-time, quantum critical behavior arises in the periodic Ising spin chain when the imaginary longitudinal magnetic field is tuned to its critical value from below. Even though the Hamiltonian is not Hermitian, the energy levels are real due to the $\mathcal{PT}$ symmetry. In this paper, we explore the analogous quantum critical behavior in higher dimensional non-Hermitian Hamiltonians on regularized spheres $S^{d-1}$. For $d=3$, we use the recently invented, powerful fuzzy sphere method, as well as discretization by the platonic solids cube, icosahedron and dodecaherdron. The low-lying energy levels and structure constants we find are in agreement with expectations from the conformal symmetry. The energy levels are in good quantitative agreement with the high-temperature expansions and with Padè extrapolations of the $6-\epsilon$ expansions in Fisher’s $i\phi^{3}$ Euclidean field theory for the Yang-Lee criticality. In the course of this work, we clarify some aspects of matching between operators in this field theory and quasiprimary fields in the $M(2,5)$ minimal model. For $d=4$, we obtain new results by replacing the $S^{3}$ with the self-dual polytope called the $24$-cell.
@@ 1 Introduction and Summary
The Ising model is the most fundamental and widely explored statistical model, which has a broad range of applications across sciences. Its one-dimensional version was solved by Ernst Ising in 1925 and shown not to possess a finite-temperature phase transition [1]. In all dimensions $d>1$, however, it has a line of first-order transitions $T<T_{c}$, which terminates at the critical point where the $\mathbb{Z}_{2}$ symmetry is restored. The properties of this second-order phase transition have been studied using many different methods, analytical and numerical. For $d>4$, the critical exponents of the Ising model are given exactly by Landau’s mean field theory, but for $d<4$ this is no longer the case. Determining the dimension dependence of the critical exponents is a fascinating problem, which continues to attract a great deal of attention.
@@
In $d=2$, the exact Ising critical exponents have been known for many years thanks to the Onsager solution [2]. The critical 2D Ising model corresponds to the simplest unitary Belavin-Polyakov-Zamolodchikov (BPZ) conformal minimal model $M(3,4)$ [3]. More generally, the critical $d$-dimensional Ising model is described by the continuum theory of a massless scalar field $\phi$ with the interaction term $\sim\phi^{4}$. The problem of finding the critical exponents is then mapped to calculation of the scaling dimensions of operators $\phi$, $\phi^{2}$, $\phi^{4}$, etc. In dimensions slightly below $4$, the IR fixed point is weakly coupled, and the Wilson-Fisher $4-\epsilon$ expansion [4] provides a useful method for approximating the $d$ dependence of operator dimensions. This has led to estimates of the 3D Ising critical exponents that are in very good agreement with experiments and Monte Carlo simulations. A significant improvement in accuracy was provided by the numerical conformal bootstrap [5, 6], which is based on the conformal symmetry of critical fluctuations [7] (for reviews, see [8, 9, 10, 11]). Furthermore, the numerical bootstrap method has been applied to the operator scaling dimensions in the entire range $42\leq d<$ 4 [12, 13], producing excellent agreement with resummed $4-\epsilon$ expansions for the $\phi^{4}$ theory.
@@
More recently, new methods for studying the 3D critical phenomena have emerged [14, 15, 16, 17, 18]. They rely on the critical behavior of the quantum theory in $2+1$ dimensions, which is analogous to the well-known periodic quantum spin chains in $1+1$ dimensions. In the continuum limit, the latter give the energy levels on a circle which, via the 2D Conformal Field Theory state-operator map determine the scaling dimensions of the primary operators and their descendants. To study a 3D CFT in a similar way, one must define the quantum Hamiltonian system on an appropriately regularized sphere $S^{2}$. In an important development [14, 15, 19], the fuzzy sphere regularization [20, 21] was shown to work very nicely for the critical 3D Ising model, giving the low-energy states that organize into the expected conformal multiplets (in fact, replacing the $S^{2}$ by an icosahedron works quite well too [17], even though this method does not readily allow for extrapolation to the continuum limit). The fuzzy sphere measurements of various CFT data approximately satisfy the conformal symmetry; this provides non-trivial checks and also can be used to correct the finite-size errors in this approach [18]. The conformal symmetry generators can be directly constructed from the fuzzy sphere observables [22, 23]. The fuzzy sphere approach has also been used to study various other 3D CFTs and defect CFTs[24, 25, 26, 27, 28, 29, 30].
Besides the scaling dimensions and OPE coefficients in a 3D CFT, the fuzzy sphere approach provides access to the sphere free energy $F$, which characterizes the number of degrees of freedom [31, 32, 33]. This quantity can be extracted from the entanglement entropy across the equator of $S^{2}$, and the numerical result for $F$ in the Ising model [16] is in excellent agreement with the extrapolation of the $4-\epsilon$ expansion of $F$ [34].
@@
A fascinating feature of the $d$-dimensional Ising model is the appearance of zeros of the partition function $Z$ at imaginary values of the magnetic field [35, 36]. In the paramagnetic phase $T>T_{c}$, these Yang-Lee (YL) zeros become dense near the edge value $h_{\rm crit}(T)$. As shown by Kortman and Griffiths [37], the density of zeros has the scaling form $\rho(h)\sim(h-h_{\rm crit}(T))^{\sigma}$, and $\sigma$ is negative in $d=2$ but positive in $d=3$. In an important further development, Michael Fisher [38] realized that this critical behavior is described by the Euclidean field theory with the imaginary potential $\sim i\phi^{3}$. In this continuum approach, the critical exponent is $\sigma=\frac{\Delta_{\phi}}{d-\Delta_{\phi}}$, where $\Delta_{\phi}$ is the scaling dimension of the real scalar field $\phi$. The field theory description of the YL universality class implies that its upper critical dimension is $6$. The lower critical dimension is $1$, and the exact solution there gives $\sigma=-\frac{1}{2}$ [35, 36]. Therefore, an interesting problem is to determine the $d$ dependence of the critical exponents in the range $61<d<$ 6. One approach, analogous to the Wilson-Fisher $4-\epsilon$ expansion for the Ising universality class, is to develop the $6-\epsilon$ expansion for the massless $i\phi^{3}$ field theory. This approach, in conjunction with the exact result in $d=1$, was used by Fisher [38] to estimate the $d$-dependence of $\sigma$ for the YL universality class. Soon after the advent of the BPZ minimal models [3], Cardy [39] realized that the critical YL model is also exactly solvable in $d=2$, where it corresponds to the simplest non-unitary minimal model $M(2,5)$ (for a recent review, see [40]). The field $\phi$ corresponds to the Virasoro primary $\Phi_{1,2}$ of holomorphic dimension $h_{1,2}=-\frac{1}{5}$. This means that $\Delta_{\phi}=-\frac{2}{5}$ and $\sigma=-\frac{1}{6}$ in $d=2$.
@@
In the range $62<d<$ 6, no similar exact results are known. In this paper, we will use a combination of $6-\epsilon$ expansions and numerical diagonalizations to approximately determine the scaling dimensions of various operators. The $6-\epsilon$ expansions are now available to $O(\epsilon^{5})$ [41]. Their resummations have been extrapolated to $d=2$, yielding good agreement with $M(2,5)$. However, for the higher dimension operators this method can exhibit significant uncertainties. The numerical boostrap approach is complicated by the fact that the YL model is non-unitary, so that the standard methods based on positivity cannot be used. A different bootstrap approach to the YL model was used in [42, 43, 44], but it does not appear to be as precise as the numerical bootstrap based on positivity. We will instead follow the successes of the fuzzy sphere [14] and the icosahedron[17] [^1] The icosahedron has also been used to discretize the 2D classical Ising model on a sphere [45, 46] approaches to the 3D Ising model, adapting them to the YL case where a non-Hermitian Hamiltonian has to be tuned to quantum criticality. We will proceed analogously to the quantum spin chain approach which has proven to be successful for describing the 2D YL model. Here an imaginary longitudinal magnetic field $h_{z}$ was introduced and tuned to a critical value from below [47, 48, 49]. Even though the spin chain Hamiltonian is complex, the energy levels are real for $h_{z}<h_{z}^{\rm crit}$ due to the $\mathcal{PT}$ symmetry [49].
@@
This paper is organized as follows. In section 2 we review the minimal model $M(2,5)$ and its field theory description, discussing the correspondence between the two. In section 3 we review the quantum spin chain with an imaginary longitudinal magnetic field [50, 47, 48, 49]. We carry out its numerical diagonalizations and show that the appropriate criteria for quantum criticality, analogous to those used recently [51] for a precise study of the Ising transition in the Schwinger model, lead to excellent agreement with the exact results from the $M(2,5)$ minimal model. In section 4 we study a quantum spin model with an imaginary longitudinal magnetic field, putting it on the platonic solids cube, icosahedron and dodecahedron, which are viewed as regularized versions of $S^{2}$. We obtain good agreement with the expected YL criticality. In section 5, we carry out analogous calculations on the fuzzy sphere and extrapolate them to the thermodynamic limit. Our fuzzy sphere result for the scaling dimension of $\phi$, $\Delta_{\phi}=0.214(2)$ is in excellent agreement with the high-temperature expansion result $0.214(6)$ from [52] and is also very close to our two-sided Padé estimate $\approx 0.218$. In section 6, we obtain first results for the quantum criticality of the $d=4$ YL model by replacing the $S^{3}$ by the $24$-cell, which is a selfdual regular $4$-polytope.
@@ 2 Minimal model $M(2,5)$ and its field theory description
The lowest level null vectors in the $M(2,5)$ model are [56, 57]
$L_{-1}|\textrm{I}\rangle,\quad\Big{(}L_{-2}^{2}+\frac{3}{5}L_{-4}\Big{)}|\textrm{I}\rangle,\quad\Big{(}L_{-2}-\frac{5}{2}L_{-1}^{2}\Big{)}|\phi\rangle,\quad\Big{(}L_{-3}-\frac{10}{3}L_{-1}L_{-2}+\frac{25}{36}L_{-1}^{3}\Big{)}|\phi\rangle\,.$ (2.3)
All the operators (primary and descendants) up to level $4$ are exhibited in Figure 1, where $T=L_{-2}|\textrm{I}\rangle$ is the holomorphic part of the stress-energy tensor and
$Q\phi=(L_{-4}-\frac{625}{624}L_{-1}^{4})|\phi\rangle\ ,\qquad\bar{Q}\phi=(\bar{L}_{-4}-\frac{625}{624}\bar{L}_{-1}^{4})|\phi\rangle\ .$ (2.4)
denote the spin $4$ quasi-primary state [56]. A related spin 00 quasiprimary is $\Xi=Q\bar{Q}|\phi\rangle$ [57].
Using the operator $Q_{6}=L_{-6}-\frac{35}{43}L_{-5}L_{-1}-\frac{105}{43}L_{-4}L_{-2}$, we can also construct a spin $6$ quasiprimary state [58, 59]:
$Q_{6}|\phi\rangle\ ,\qquad\bar{Q}_{6}|\phi\rangle\ \ .$ (2.5)
@@
The operator $\phi^{2}$ is a descendant by the equation of motion. Therefore, the next spin-zero conformal primary field is $\phi^{3}$. Its scaling dimension is [61, 62, 63, 41]
@@
The field theory also contains primary operators with an even number of derivatives, which have schematic form $\phi\partial_{\mu_{1}}\partial_{\mu_{2}}\ldots\partial_{\mu_{2n}}\phi+\ldots$. Projecting onto the traceless part of such an operator, we find the leading twist operators of spin $2n$ (they form the leading Regge trajectory). For $n=1$ this operator is the conserved stress-energy tensor, but for $n>1$ they are not conserved. The first such operator is the traceless rank $4$ tensor
@@
Table 2: Two-sided Padé approximations of the scaling dimensions of lowest primary operators in $d=3,4,5$. All of them are below the unitarity bounds.
@@ 3 Spin chain for 2D Yang-Lee model
Figure 4: The qualitative phase diagram of the quantum Ising model with transverse magnetic field $h_{x}$ and imaginary longitudinal magnetic field $h_{z}$. The $h_{z}=0$ horizontal line is the $\mathbb{Z}_{2}$ symmetric line. The Ising critical point, which is located at this line, is shown by the green dot. To the right of this point, the Ising model is in the paramagnetic phase, also known as the high temperature phase, shown by the black section of the $\mathbb{Z}_{2}$ symmetric line. The paramagnetic phase with a small $h_{z}$ deformation still preserves $\mathcal{PT}$ symmetry, shown by the white region. When $h_{z}$ passes its critical value, the $\mathcal{PT}$ symmetry is spontaneously broken, shown by the blue region. The critical coupling $h_{z}$ depends on $h_{x}$ and therefore we have a line of critical points corresponding to the YL criticality, shown as the blue solid line. We study the Ising to YL criticality flow by moving along the red dashed arrows between points a, b and c.
@@
In the numerical spectrum, this ratio can be written using the energy levels of distinct spin sectors, namely, $E^{(\textbf{s})}_{n}$ with $n=0,1,2,\dots$ represents the $n$th energy level in spin sector $\textbf{s}=0,1,2,\dots$, so
@@
to be exactly $r_{T}=3$. Numerically this ratio can be obtained using the following energy levels in different irreps of $O$ and $I$:
@@
where we used that $\textbf{5}_{O(3)}$ splits as $\textbf{2}+\textbf{3}^{\prime}$ under the cubic symmetry group $O$, whereas it does not split for Icosahedral group $I$, and $\textbf{1}_{O(3)}$ and $\textbf{3}_{O(3)}$ do not split for $O$ and $I$ and corresponds to irrep 1 and 3. Here we list values $ih_{z}^{*}$, where $r_{T}=3$ in the Table 7. They are slightly away from the merger points listed in the previous Table 6.
@@
$\textrm{Cube}:\quad\Delta_{\phi}=3\frac{E_{1}^{(\textbf{1})}-E_{0}^{(\textbf{1})}}{E_{1}^{(\textbf{3}^{\prime})}-E_{0}^{(\textbf{1})}}\,,\qquad\textrm{Icosahedron/Dodecahedron}:\quad\Delta_{\phi}=3\frac{E_{1}^{(\textbf{1})}-E_{0}^{(\textbf{1})}}{E_{1}^{(\textbf{5})}-E_{0}^{(\textbf{1})}}\,,$ (4.4)